of the $50,000 that Natasha pocketed on her last real estate deal, $20,000 went to charity. She invested part of the remainder with an annuak yield of 16% and the rest with an annual yield of 25%. if she made $6060.00 on these investments in one year, then how much did she invest in each fund?
Answer by stanbon(48535)

of the $50,000 that Natasha pocketed on her last real estate deal, $20,000 went to charity. She invested part of the remainder with an annuak yield of 16% and the rest with an annual yield of 25%. if she made $6060.00 on these investments in one year, then how much did she invest in each fund?
-----------------------
16% DATA:
Amt invested = x dollars; interest = 0.16x dollars
---------------
25% DATA:
Amt invested = 30,000-x dollars; interest = 0.25(30,000-x)=7500-0.25x dollars
--------------
EQUATION:
interest + interest = 6060 dollars
0.16x + 7500 - 0.25x = 6060
-0.09x = -1440
x = 16000 (amt invesed at 16%)
30000-16000= 14000 (amt invested at 25%)
==========================
Cheers,
Stan H.